# Governance Architecture for Autonomous Agent Systems: Threats, Framework, and Engineering Practice

### 07 Mar 2026 v1

@@ Abstract.
Autonomous agents powered by large language models introduce a class of execution-layer vulnerabilities—prompt injection, retrieval poisoning, and uncontrolled tool invocation—that existing guardrails fail to address systematically. In this work, we propose the Layered Governance Architecture (LGA), a four-layer framework comprising execution sandboxing (L1), intent verification (L2), zero-trust inter-agent authorization (L3), and immutable audit logging (L4). To evaluate LGA, we construct a bilingual benchmark (Chinese original, English via machine translation) of 1,081 tool-call samples—covering prompt injection, RAG poisoning, and malicious skill plugins—and apply it to OpenClaw, a representative open-source agent framework. Experimental results on Layer 2 intent verification with four local LLM judges (Qwen3.5-4B, Llama-3.1-8B, Qwen3.5-9B, Qwen2.5-14B) and one cloud judge (GPT-4o-mini) show that all five LLM judges intercept 93.0–98.5% of TC1/TC2 malicious tool calls, while lightweight NLI baselines remain below 10%. TC3 (malicious skill plugins) proves harder at 75–94% IR among judges with meaningful precision–recall balance, motivating complementary enforcement at Layers 1 and 3. Qwen2.5-14B achieves the best local balance (98% IR, 10–20% FPR); a two-stage cascade (Qwen3.5-9BGPT-4o-mini) achieves 91.9–92.6% IR with 1.9–6.7% FPR; a fully local cascade (Qwen3.5-9BQwen2.5-14B) achieves 94.7–95.6% IR with 6.0–9.7% FPR for data-sovereign deployments. An end-to-end pipeline evaluation () demonstrates that all four layers operate in concert with 96% IR and a total P50 latency of 980 ms, of which the non-judge layers contribute only 18 ms. Generalization to the external InjecAgent benchmark yields 99–100% interception, confirming robustness beyond our synthetic data.
@@ 1. Introduction
LLM-driven agent systems are undergoing a paradigm shift from conversational to executive: frameworks such as AutoGen (Wu et al., 2024), LangChain (Chase, 2022), and OpenClaw111OpenClaw: https://github.com/openclaw/openclaw; see Sect. 6 for a detailed case study. now grant models the ability to write files, execute shell commands, and invoke transactional APIs. This shift extends the blast radius of failure from the cognitive layer (incorrect text) to the execution layer (irreversible system state changes).
Concurrently, AI-assisted programming tools (Chen et al., 2021) are reshaping the value distribution of software engineers. A common argument holds that low-level systems expertise is amplified by AI because models generate code riddled with leaky abstractions (Spolsky, 2002)—subtle semantic failures that only surface under production conditions. While empirically supported in the short term, we argue this framing overlooks a convergence trend: multi-agent adversarial verification, formal-methods assistance, and execution-feedback loops are systematically reducing AI defect rates, shifting the locus of engineering effort from defect remediation toward system governance. This governance imperative is sharpened by the defining property of LLM-based agents: they convert natural-language instructions into executable tool calls without human mediation. LGA preserves this autonomous execution for verified calls; human review is required only for escalated blocks, not the normal execution path.
Existing defenses operate at the wrong layer. Content-safety systems such as Llama Guard (Inan et al., 2023) and NeMo Guardrails (Rebedea et al., 2023) filter harmful text generation but cannot intercept unauthorized tool invocations that arise from semantically benign-looking inputs. Agent frameworks such as LangChain (Chase, 2022) and AutoGen (Wu et al., 2024) provide tool-chaining interfaces but lack execution-layer isolation. Benchmarks such as InjecAgent (Zhan et al., 2024) and AgentDojo (Debenedetti et al., 2024) quantify attack success rates but do not propose deployable mitigations. No existing work addresses all three threat classes—prompt injection, RAG poisoning, and malicious skill plugins—within a unified, independently deployable governance architecture.
@@
We identify a systematic gap in existing agent defenses—they operate at the text-generation layer and cannot intercept execution-layer threats—and argue that the resulting governance imperative motivates a shift from defect remediation toward system-level invariants (Sect. 3).
@@ 2.1. LLM Code Generation and Abstraction Failures
Spolsky’s Law of Leaky Abstractions (Spolsky, 2002) posits that non-trivial abstractions inevitably expose implementation details. Chen et al. (Chen et al., 2021) demonstrate that Codex (2021) achieves 28.8% pass@1 on HumanEval—a figure that has since improved dramatically, yet the underlying pattern persists: LLMs degrade significantly on tasks involving low-level concurrency semantics. Liu et al. (Liu et al., 2023a) show that out-of-distribution (OOD) code generated by LLMs exhibits a substantially higher security-vulnerability rate than in-distribution code. These findings support the short-term value of systems expertise but do not address its long-term trajectory. For agent systems, execution-layer vulnerabilities—such as unverified tool calls that cause irreversible state changes—persist regardless of code quality, motivating the governance-first approach of LGA.
@@ 2.3. Agent Frameworks
AutoGPT (Significant Gravitas, 2023) and AutoGen (Wu et al., 2024) represent two dominant design philosophies: single-agent autonomous loops versus multi-agent orchestration. LangChain (Chase, 2022) provides standardized tool-chaining interfaces but relies on input filtering at the application layer, lacking physical execution-layer isolation. Existing frameworks leave a systematic gap in governance architecture, which LGA addresses.
@@ 3. From Defect Remediation to System Governance
Crucially, these governance boundaries are non-negotiable invariants: no amount of model capability improvement can substitute for execution-layer isolation or cryptographic authorization, because their correctness is informed by legal and contractual requirements and enforced through architectural mechanisms rather than statistical generalization. Sect. 5 operationalizes this principle into a concrete four-layer architecture.
@@ 4.2. Threat Class 1: Agency Abuse via Prompt Injection
In agents with physical execution rights, a successful Agency Abuse attack can exfiltrate filesystem contents, issue unauthorized network requests, or modify database state—all without human intervention and within milliseconds of receiving the adversarial input.
@@ 4.3. Threat Class 2: RAG Data Poisoning
In multi-agent networks, an infected agent may continue writing poisoned entries during task execution, forming a positive-feedback infection loop with potentially cascading propagation.
@@ 5.1. Layer 1: Execution Sandbox
All agent code execution is confined within OS-level isolation. A full production deployment would employ Linux containers with seccomp syscall filtering, read-only filesystem mounts, and path-whitelist enforcement; our experimental prototype uses the lighter-weight unshare namespace isolation (Sect. 7.5). The critical design principle is that isolation must be implemented at the OS layer, not the software layer: software-layer access controls can be bypassed by sufficiently crafted prompt injections, whereas OS-level physical isolation is orthogonal to agent semantics.
@@ 6.1. Architecture Overview
We select OpenClaw222OpenClaw is an open-source local agent framework. Analysis based on the publicly available source code at https://github.com/openclaw/openclaw (commit f014e25, accessed 2026-03-05). as our case study because it implements all four architectural components that LGA targets—local persistent memory, an autonomous execution loop, a compute-routing gateway, and a plugin ecosystem—providing a concrete testbed for compliance gap analysis. OpenClaw is a local autonomous agent system whose core design philosophy is to “enclose AI capabilities within governable physical boundaries” (OpenClaw Contributors, 2025). Its principal technical components are: Zero-Ops local persistent memory, a heartbeat autonomous loop with SOUL.md configuration, a model-selection gateway, and a community skill ecosystem (ClawHub).
@@ 7.4. Discussion
The results reveal a stark capability gap between lightweight NLI models and LLM-based judges. BART-MNLI achieves only 8–10% interception despite zero false positives and a moderate median overhead of 204–258 ms. All five LLM judges (Qwen3.5-4B through GPT-4o-mini) achieve at least 93% overall interception on TC1/TC2 (across all 457 malicious samples; individual attack-type rates may be slightly lower), but with markedly different trade-off profiles. Among local models, Qwen2.5-14B achieves the best overall balance (98.2–98.5% IR, 9.7–20.1% FPR, 880–1,160 ms), approaching GPT-4o-mini’s precision at local-deployment cost. Llama-3.1-8B and Qwen3.5-9B exhibit a high-recall, low-precision profile (95% IR) but at elevated FPR (23.7–37.7%), suggesting a conservative blocking bias. Qwen3.5-4B, the smallest model (3.4 GB), still achieves 94.3–95.8% IR with the lowest latency among LLM judges (490 ms), making it viable for latency-sensitive edge deployments. GPT-4o-mini (93.0–95.4% IR, 3.2–12.3% FPR, 1,100 ms) remains the most precise option but requires cloud connectivity. The two cascade conditions (2,800–3,000 ms) occupy the highest-latency, lowest-FPR region of the trade-off space, offering the best precision at the cost of throughput. The NLI models’ near-zero interception indicates that zero-shot textual entailment is insufficient for tool-call authorization—the task requires structured reasoning about intent–action alignment, which current NLI architectures cannot capture without task-specific fine-tuning. ROC analysis on TC1/TC2 ( per language: 457 malicious + 464 non-plugin benign; the 40 plugin hard negatives are excluded to match the FPR computation scope) confirms this gap: using each NLI model’s continuous entailment score as the ranking function (converted to a malicious-class score via , applied uniformly regardless of binary verdict), BART-MNLI achieves AUC = 0.75 (ZH) / 0.88 (EN)—the EN AUC indicates reasonable ranking, yet the optimal binary threshold still yields only 8.1% IR because both classes’ entailment scores overlap substantially in the decision-relevant range—and mDeBERTa-NLI AUC 0.50 (ZH) / 0.56 (EN; continuous scores only, no binary threshold evaluation; see Table 3 footnote), near random; by contrast, the five LLM judges—which produce binary ALLOW/BLOCK verdicts—achieve single operating points in the upper-left region of ROC space (TPR 0.93, FPR 0.36), well above both NLI curves at every threshold.
@@ 7.4.2. Attack Subtype Variation
The representative LLM judges (GPT-4o-mini and Qwen3.5-9B) show high interception across all attack subtypes. GPT-4o-mini exhibits remarkably uniform rates: 92.9% (direct), 92.8% (indirect), and 93.4% (RAG-poisoning). Qwen3.5-9B achieves even higher rates on direct injection (97.2%) and RAG-poisoning (98.5%), but slightly lower on indirect injection (94.4%). Interestingly, BART-MNLI shows the opposite pattern: RAG-poisoning attacks are intercepted at 17.6% versus only 4.3% for direct injection; this pattern holds in English (11.8% vs. 4.3%) though the gap is smaller. This suggests that RAG-poisoning attacks produce more semantically anomalous tool calls that even shallow NLI models can partially detect, while direct injections craft more plausible-looking tool calls. Notably, Qwen2.5-14B achieves perfect direct-injection interception (100% D on both ZH and EN), the only perfect subtype score in Table 3; this likely reflects the high semantic anomaly of direct injections for larger models rather than a data artifact.
@@ 7.4.3. Threat Class 3: Malicious Skill Plugins
The Plugin column (P) in Table 3 reveals a markedly lower interception rate compared to TC1/TC2 subtypes: GPT-4o-mini achieves 75.0% (ZH) and 94.2% (EN), Qwen2.5-14B reaches 77.5% (ZH) and 88.3% (EN), and Qwen3.5-9B reaches 77.5% on both languages—versus 92–100% on direct/indirect injection and RAG poisoning. BART-MNLI blocks everything on Chinese (100% IR / 100% FPR) yet reaches only 63.3% on English, confirming that NLI cannot reliably detect permission-boundary violations. This gap is expected: plugin attacks perform their declared operation while adding a covert side-channel (e.g., telemetry exfiltration), so the tool call is partially consistent with the task, making the violation harder to detect from intent–action alignment alone.
Table 6 reports the per-subtype breakdown ( per subtype). Privilege escalation is the most reliably detected (86.7–93.3%), likely because system-level operations (file permission changes, admin API calls) diverge sharply from declared plugin scope. Supply-chain attacks are also well-detected (83.3–96.7%), as references to external downloads or out-of-scope URLs are unambiguously anomalous. Exfiltration and persistence are harder (46.7–96.7%), as their malicious parameters blend into legitimate-looking telemetry or configuration fields. Per-subtype rates vary across languages and models (e.g., GPT-4o-mini EN achieves 90.0–96.7% across all subtypes, versus 46.7–90.0% on ZH), reinforcing the cross-lingual asymmetry discussed below. The persistence subtype shows the most extreme divergence for GPT-4o-mini (ZH: 46.7%, EN: 96.7%, 50 pp gap), consistent with reduced Chinese-language sensitivity to persistence-related semantic patterns (e.g., scheduled execution, startup registration) rather than a data artifact, as Qwen3.5-9B shows a much smaller gap on the same subtype (60.0% vs. 70.0%). These results motivate complementary defenses at Layer 1 (sandboxing prevents actual exfiltration) and Layer 3 (declared-permission enforcement at the protocol level) to compensate for Layer 2’s reduced detection rate on TC3.
@@
A critical finding is the divergence between Qwen3.5-9B’s high interception rate and its elevated false positive rate (34.1% on Chinese, 23.7% on English). This conservatism likely reflects the model’s safety alignment: Qwen3.5-9B tends to block ambiguous tool calls as a precaution, whereas GPT-4o-mini shows better calibration between security and usability. In production systems, this trade-off can be addressed through LGA’s layered design: Layer 1 (sandbox) provides physical containment for false negatives, while Layer 2’s false positives are recoverable via human-in-the-loop escalation or cascaded verification. Notably, Qwen3.5-9B’s high interception rate—achieved without any same-source advantage over the benchmark—confirms that the security capability of LLM judges is genuine and not an artifact of data contamination; the elevated FPR is a calibration issue, not a fundamental capability limitation. However, with 24–34% FPR, standalone Qwen3.5-9B is impractical for autonomous deployment—at a low attack prevalence (), its positive predictive value drops below 4% (2.8% ZH / 3.9% EN; see Table 7). Qwen3.5-9B is therefore best deployed as a high-recall first stage in a cascade or with human-in-the-loop escalation, rather than as a standalone gatekeeper. For data-sovereign scenarios requiring a single local judge, Qwen2.5-14B offers a substantially better alternative (98.2–98.5% IR, 9.7–20.1% FPR depending on language).
@@
To validate the cascading mitigation, we simulate a two-stage pipeline using existing per-sample predictions: Qwen3.5-9B serves as the first-stage filter, and calls it blocks are re-evaluated by GPT-4o-mini; the final verdict is BLOCK if and only if GPT-4o-mini also issues BLOCK. On the Chinese dataset, cascading reduces FPR from 34.1% to 1.9% while retaining 91.9% IR (vs. 96.5% single-stage); on English, FPR drops from 23.7% to 6.7% with 92.6% IR (vs. 95.0%). Note that these results are based on simulated cascading (post-hoc composition of independent per-sample predictions), not live two-stage inference; actual cascade latency and decision coupling may differ slightly. Nevertheless, the cascaded scores—94.9% (Chinese) and 92.9% (English)—match or slightly exceed single-model GPT-4o-mini (94.9% and 91.4% respectively; the ZH values are equal after rounding), confirming that cascading effectively recovers usability without sacrificing the security floor established by the local model.
@@
For fully data-sovereign deployments, we also evaluate a local-only cascade (Qwen3.5-9BQwen2.5-14B) using the same post-hoc composition method. On Chinese, this achieves 95.6% IR with 6.0% FPR (); on English, 94.7% IR with 9.7% FPR (). Compared to the cloud cascade, the local variant trades a higher FPR (6.0% vs. 1.9% ZH; 9.7% vs. 6.7% EN) for higher IR (95.6% vs. 91.9% ZH; 94.7% vs. 92.6% EN), while achieving similar (within 0.1–0.4 pp); note that IR differences between the two cascades are statistically significant (Table 8), but is not separately tested as it compounds IR and FPR in opposite directions. This confirms that a viable local-only cascade exists for scenarios where cloud APIs are prohibited.

### 10 Mar 2026 v2

@@ Abstract.
Autonomous agents powered by large language models introduce a class of execution-layer vulnerabilities—prompt injection, retrieval poisoning, and uncontrolled tool invocation—that existing guardrails evaluated in our study do not address within a unified architecture. In this work, we propose the Layered Governance Architecture (LGA), a four-layer framework comprising execution sandboxing (L1), intent verification (L2), zero-trust inter-agent authorization (L3), and immutable audit logging (L4). To evaluate LGA, we construct a bilingual benchmark (Chinese original, English via machine translation) of 1,081 tool-call samples—covering prompt injection, RAG poisoning, and malicious skill plugins—and apply it to OpenClaw, a representative open-source agent framework. Experimental results on Layer 2 intent verification with four local LLM judges (Qwen3.5-4B, Llama-3.1-8B, Qwen3.5-9B, Qwen2.5-14B) and one cloud judge (GPT-4o-mini) show that all five LLM judges intercept 93.0–98.5% of TC1/TC2 malicious tool calls, while lightweight NLI baselines remain below 10%. TC3 (malicious skill plugins) proves harder at 75–94% IR among judges with meaningful precision–recall balance, motivating complementary enforcement at Layers 1 and 3. Qwen2.5-14B achieves the best local balance (98% IR, 10–20% FPR); a two-stage cascade (Qwen3.5-9BGPT-4o-mini) achieves 91.9–92.6% IR with 1.9–6.7% FPR; a fully local cascade (Qwen3.5-9BQwen2.5-14B) achieves 94.7–95.6% IR with 6.0–9.7% FPR for data-sovereign deployments. An end-to-end pipeline evaluation () demonstrates that all four layers operate in concert with 96% IR and a total P50 latency of 980 ms, of which the non-judge layers contribute only 18 ms. Generalization to the external InjecAgent benchmark yields 99–100% interception, providing evidence of generalization beyond our synthetic data.
@@ 1. Introduction
LLM-driven agent systems are transitioning from conversational to executive: frameworks such as AutoGen (Wu et al., 2024), LangChain (Chase, 2022), and OpenClaw111OpenClaw: https://github.com/openclaw/openclaw; see Sect. 6 for a detailed case study. now grant models the ability to write files, execute shell commands, and invoke transactional APIs. This shift moves the consequence of failure from incorrect text output to potentially irreversible system state changes.
Concurrently, AI-assisted programming tools (Chen et al., 2021) are reshaping the value distribution of software engineers. A common argument holds that low-level systems expertise is amplified by AI because models generate code containing leaky abstractions (Spolsky, 2002)—semantic failures that may surface under production conditions. While empirically supported in the short term, we argue this framing overlooks a convergence trend: multi-agent adversarial verification, formal-methods assistance, and execution-feedback loops are systematically reducing AI defect rates, shifting the locus of engineering effort from defect remediation toward system governance. This governance concern is sharpened by the defining property of LLM-based agents: they convert natural-language instructions into executable tool calls without human mediation. LGA preserves this autonomous execution for verified calls; human review is required only for escalated blocks, not the normal execution path.
Existing defenses target the text-generation layer and do not intercept unauthorized tool invocations arising from semantically benign-looking inputs. Content-safety systems such as Llama Guard (Inan et al., 2023) and NeMo Guardrails (Rebedea et al., 2023) filter harmful text generation but lack tool-call–level interception (see Sect. 2 for a survey). Agent frameworks such as LangChain (Chase, 2022) and AutoGen (Wu et al., 2024) provide tool-chaining interfaces but lack execution-layer isolation. Benchmarks such as InjecAgent (Zhan et al., 2024) and AgentDojo (Debenedetti et al., 2024) quantify attack success rates but do not propose deployable mitigations. To our knowledge, no existing work simultaneously addresses all three threat classes—prompt injection, RAG poisoning, and malicious skill plugins—within a unified governance architecture; Sect. 2 surveys the closest prior art.
@@
We identify a systematic gap in existing agent defenses—they operate at the text-generation layer and cannot intercept execution-layer threats—and argue that the resulting governance concern motivates a shift from defect remediation toward system-level invariants (Sect. 3).
@@ 2.1. LLM Code Generation and Abstraction Failures
Spolsky’s Law of Leaky Abstractions (Spolsky, 2002) posits that non-trivial abstractions inevitably expose implementation details. Chen et al. (Chen et al., 2021) demonstrate that Codex (2021) achieves 28.8% pass@1 on HumanEval—a figure that has since improved substantially, yet the underlying pattern persists: LLMs degrade significantly on tasks involving low-level concurrency semantics. Liu et al. (Liu et al., 2023a) show that out-of-distribution (OOD) code generated by LLMs exhibits a substantially higher security-vulnerability rate than in-distribution code. These findings support the short-term value of systems expertise but do not address its long-term trajectory. For agent systems, execution-layer vulnerabilities—such as unverified tool calls that cause irreversible state changes—persist regardless of code quality, motivating the governance-first approach of LGA.
@@ 2.3. Agent Frameworks
AutoGPT (Significant Gravitas, 2023) and AutoGen (Wu et al., 2024) represent two dominant design philosophies: single-agent autonomous loops versus multi-agent orchestration. LangChain (Chase, 2022) provides standardized tool-chaining interfaces but relies on input filtering at the application layer, lacking physical execution-layer isolation. These frameworks do not provide a unified governance architecture, a gap that LGA aims to address.
@@ 3. From Defect Remediation to System Governance
These governance boundaries are enforced through architectural mechanisms rather than statistical generalization, and cannot be relaxed without explicit policy changes: current model capability improvements do not substitute for execution-layer isolation or cryptographic authorization, because their correctness derives from legal and contractual requirements rather than optimizable objectives (e.g., at 1% attack prevalence even the best local judge achieves only 22.7% PPV; see Table 7). Sect. 5 operationalizes this principle into a concrete four-layer architecture.
@@ 4.2. Threat Class 1: Agency Abuse via Prompt Injection
In agents with physical execution rights, a successful Agency Abuse attack can exfiltrate filesystem contents, issue unauthorized network requests, or modify database state—potentially without human intervention before irreversible state changes occur.
@@ 4.3. Threat Class 2: RAG Data Poisoning
In multi-agent networks, an infected agent may continue writing poisoned entries during task execution, forming a positive-feedback loop with potentially cascading propagation across agents.
@@ 5.1. Layer 1: Execution Sandbox
All agent code execution is confined within OS-level isolation. A full production deployment would employ Linux containers with seccomp syscall filtering, read-only filesystem mounts, and path-whitelist enforcement; our experimental prototype uses the lighter-weight unshare namespace isolation (Sect. 7.5). The key design principle is that isolation should be implemented at the OS layer rather than the software layer: software-layer access controls are subject to semantic manipulation (as demonstrated in (Zhan et al., 2024)), whereas OS-level isolation operates independently of the agent’s language processing.
@@ 6.1. Architecture Overview
We select OpenClaw222OpenClaw is an open-source local agent framework. Analysis based on the publicly available source code at https://github.com/openclaw/openclaw (commit f014e25, accessed 2026-03-05). as our case study because it implements all four architectural components that LGA targets—local persistent memory, an autonomous execution loop, a compute-routing gateway, and a plugin ecosystem—providing a concrete testbed for compliance gap analysis. OpenClaw is a local autonomous agent system whose stated design goal is to “enclose AI capabilities within governable physical boundaries” (OpenClaw Contributors, 2025); we evaluate whether the implementation achieves this goal. Its principal technical components are: Zero-Ops local persistent memory, a heartbeat autonomous loop with SOUL.md configuration, a model-selection gateway, and a community skill ecosystem (ClawHub).
@@ 7.4. Discussion
The results reveal a clear capability gap between lightweight NLI models and LLM-based judges. BART-MNLI achieves only 8–10% interception despite zero false positives and a moderate median overhead of 204–258 ms. All five LLM judges (Qwen3.5-4B through GPT-4o-mini) achieve at least 93% overall interception on TC1/TC2 (across all 457 malicious samples; individual attack-type rates may be slightly lower), but with markedly different trade-off profiles. Among local models, Qwen2.5-14B achieves the best overall balance (98.2–98.5% IR, 9.7–20.1% FPR, 880–1,160 ms), approaching GPT-4o-mini’s precision at local-deployment cost. Llama-3.1-8B and Qwen3.5-9B exhibit a high-recall, low-precision profile (95% IR) but at elevated FPR (23.7–37.7%), suggesting a conservative blocking bias. Qwen3.5-4B, the smallest model (3.4 GB), still achieves 94.3–95.8% IR with the lowest latency among LLM judges (490 ms), making it viable for latency-sensitive edge deployments. GPT-4o-mini (93.0–95.4% IR, 3.2–12.3% FPR, 1,100 ms) remains the most precise option but requires cloud connectivity. The two cascade conditions (2,800–3,000 ms) occupy the highest-latency, lowest-FPR region of the trade-off space, offering the best precision at the cost of throughput. The NLI models’ near-zero interception suggests that zero-shot textual entailment, as instantiated by BART-MNLI and mDeBERTa-NLI, is insufficient for tool-call authorization without task-specific adaptation. ROC analysis on TC1/TC2 ( per language: 457 malicious + 464 non-plugin benign; the 40 plugin hard negatives are excluded to match the FPR computation scope) confirms this gap: using each NLI model’s continuous entailment score as the ranking function (converted to a malicious-class score via , applied uniformly regardless of binary verdict), BART-MNLI achieves AUC = 0.75 (ZH) / 0.88 (EN)—the EN AUC indicates reasonable ranking, yet the optimal binary threshold still yields only 8.1% IR because both classes’ entailment scores overlap substantially in the decision-relevant range—and mDeBERTa-NLI AUC 0.50 (ZH) / 0.56 (EN; continuous scores only, no binary threshold evaluation; see Table 3 footnote), near random; by contrast, the five LLM judges—which produce binary ALLOW/BLOCK verdicts—achieve single operating points in the upper-left region of ROC space (TPR 0.93, FPR 0.36), well above both NLI curves at every threshold.
@@ 7.4.2. Attack Subtype Variation
The representative LLM judges (GPT-4o-mini and Qwen3.5-9B) show high interception across all attack subtypes. GPT-4o-mini exhibits relatively uniform rates: 92.9% (direct), 92.8% (indirect), and 93.4% (RAG-poisoning). Qwen3.5-9B achieves even higher rates on direct injection (97.2%) and RAG-poisoning (98.5%), but slightly lower on indirect injection (94.4%). Interestingly, BART-MNLI shows the opposite pattern: RAG-poisoning attacks are intercepted at 17.6% versus only 4.3% for direct injection; this pattern holds in English (11.8% vs. 4.3%) though the gap is smaller. This suggests that RAG-poisoning attacks produce more semantically anomalous tool calls that even shallow NLI models can partially detect, while direct injections craft more plausible-looking tool calls. Notably, Qwen2.5-14B achieves perfect direct-injection interception (100% D on both ZH and EN), the only perfect subtype score in Table 3; this likely reflects the high semantic anomaly of direct injections for larger models rather than a data artifact.
@@ 7.4.3. Threat Class 3: Malicious Skill Plugins
The Plugin column (P) in Table 3 reveals a markedly lower interception rate compared to TC1/TC2 subtypes: GPT-4o-mini achieves 75.0% (ZH) and 94.2% (EN), Qwen2.5-14B reaches 77.5% (ZH) and 88.3% (EN), and Qwen3.5-9B reaches 77.5% on both languages—versus 92–100% on direct/indirect injection and RAG poisoning. BART-MNLI blocks everything on Chinese (100% IR / 100% FPR) yet reaches only 63.3% on English, suggesting that the tested NLI models cannot reliably detect permission-boundary violations. This gap is expected: plugin attacks perform their declared operation while adding a covert side-channel (e.g., telemetry exfiltration), so the tool call is partially consistent with the task, making the violation harder to detect from intent–action alignment alone.
Table 6 reports the per-subtype breakdown ( per subtype). Privilege escalation is the most reliably detected (86.7–93.3%), likely because system-level operations (file permission changes, admin API calls) diverge sharply from declared plugin scope. Supply-chain attacks are also well-detected (83.3–96.7%), as references to external downloads or out-of-scope URLs are typically anomalous. Exfiltration and persistence are harder (46.7–96.7%), as their malicious parameters blend into legitimate-looking telemetry or configuration fields. Per-subtype rates vary across languages and models (e.g., GPT-4o-mini EN achieves 90.0–96.7% across all subtypes, versus 46.7–90.0% on ZH), reinforcing the cross-lingual asymmetry discussed below. The persistence subtype shows the most extreme divergence for GPT-4o-mini (ZH: 46.7%, EN: 96.7%, 50 pp gap), consistent with reduced Chinese-language sensitivity to persistence-related semantic patterns (e.g., scheduled execution, startup registration) rather than a data artifact, as Qwen3.5-9B shows a much smaller gap on the same subtype (60.0% vs. 70.0%). These results motivate complementary defenses at Layer 1 (sandboxing prevents actual exfiltration) and Layer 3 (declared-permission enforcement at the protocol level) to compensate for Layer 2’s reduced detection rate on TC3.
@@
A notable finding is the divergence between Qwen3.5-9B’s high interception rate and its elevated false positive rate (34.1% on Chinese, 23.7% on English). This conservatism likely reflects the model’s safety alignment: Qwen3.5-9B tends to block ambiguous tool calls as a precaution, whereas GPT-4o-mini shows better calibration between security and usability. In production systems, this trade-off can be addressed through LGA’s layered design: Layer 1 (sandbox) provides physical containment for false negatives, while Layer 2’s false positives are recoverable via human-in-the-loop escalation or cascaded verification. Notably, Qwen3.5-9B’s high interception rate—achieved without any same-source advantage over the benchmark—provides evidence that the interception capability is not solely an artifact of same-source familiarity; the elevated FPR is a calibration issue, not a fundamental capability limitation. However, with 24–34% FPR, standalone Qwen3.5-9B is impractical for autonomous deployment—at a low attack prevalence (), its positive predictive value drops below 4% (2.8% ZH / 3.9% EN; see Table 7). Qwen3.5-9B is therefore best deployed as a high-recall first stage in a cascade or with human-in-the-loop escalation, rather than as a standalone gatekeeper. For data-sovereign scenarios requiring a single local judge, Qwen2.5-14B offers a substantially better alternative (98.2–98.5% IR, 9.7–20.1% FPR depending on language).
@@
To validate the cascading mitigation, we simulate a two-stage pipeline using existing per-sample predictions: Qwen3.5-9B serves as the first-stage filter, and calls it blocks are re-evaluated by GPT-4o-mini; the final verdict is BLOCK if and only if GPT-4o-mini also issues BLOCK. On the Chinese dataset, cascading reduces FPR from 34.1% to 1.9% while retaining 91.9% IR (vs. 96.5% single-stage); on English, FPR drops from 23.7% to 6.7% with 92.6% IR (vs. 95.0%). Note that these results are based on simulated cascading (post-hoc composition of independent per-sample predictions), not live two-stage inference; actual cascade latency and decision coupling may differ slightly. Nevertheless, the cascaded scores—94.9% (Chinese) and 92.9% (English)—match or slightly exceed single-model GPT-4o-mini (94.9% and 91.4% respectively; the ZH values are equal after rounding), showing that cascading reduces FPR from 34.1% to 1.9% while retaining 91.9% IR, achieving comparable to standalone GPT-4o-mini.
@@
For fully data-sovereign deployments, we also evaluate a local-only cascade (Qwen3.5-9BQwen2.5-14B) using the same post-hoc composition method. On Chinese, this achieves 95.6% IR with 6.0% FPR (); on English, 94.7% IR with 9.7% FPR (). Compared to the cloud cascade, the local variant trades a higher FPR (6.0% vs. 1.9% ZH; 9.7% vs. 6.7% EN) for higher IR (95.6% vs. 91.9% ZH; 94.7% vs. 92.6% EN), while achieving similar (within 0.1–0.4 pp); note that IR differences between the two cascades are statistically significant (Table 8), but is not separately tested as it compounds IR and FPR in opposite directions. This indicates that a viable local-only cascade exists for scenarios where cloud APIs are prohibited.